OpenStudy (anonymous):

I need help with this. Can anyone assist me? Quadrilateral FGHI is shown on the coordinate plane with coordinates F at 4, −2; G at 7, −2; H at 5, −4; and I 2, −4. Point A is at −2, 7 and point D is at −10, −1. In order to build a scale model of the trail, the drawing is enlarged as parallelogram ABCD on the coordinate plane. If two corners of the trail are at point A (−2, 7) and point D (−10, −1), what is another point that could represent point B?

OpenStudy (iwillrektyou):

I believe you need to use the distance formula for this.

OpenStudy (anonymous):

And what exactly is that. I am clueless

OpenStudy (iwillrektyou):

Do they lie on the same line (colinear)?

OpenStudy (iwillrektyou):

And the distance formula is \[\sqrt{(x2 - x1)^2 + (y2 - y1)^2)}\]

OpenStudy (anonymous):

Yes they are colinear.

OpenStudy (iwillrektyou):

Alright. So just substitute the points for x1, x2, y1, and y2. I will guide you through this.

OpenStudy (iwillrektyou):

\[\sqrt{(-10 + 2)^2 + (-1 - 7)^2}\]

OpenStudy (anonymous):

is the answer (12, 7)

OpenStudy (iwillrektyou):

\[\sqrt{(-8)^2 + (-8)^2)}\]

OpenStudy (iwillrektyou):

\[\sqrt{64 + 64}\]

OpenStudy (iwillrektyou):

\[\sqrt{128} = \]

OpenStudy (iwillrektyou):

11.31

OpenStudy (anonymous):

Oh wow. This is really helpful. Thanks

OpenStudy (iwillrektyou):

So 11.31 is the distance between A and D. Is this quadrilateral a square or..?

OpenStudy (anonymous):

a square

OpenStudy (iwillrektyou):

Ok good! That makes it so much easier.

OpenStudy (anonymous):

So the answer options were a. (14,7) b. (10,7) c. (12,7) d, (8,7 )

OpenStudy (iwillrektyou):

Also, the points must match, so it would be A = F, B = G, C = H, D = I

OpenStudy (iwillrektyou):

\[\sqrt{(-2 - x1)^2 + (7 - y1)^2}\]

OpenStudy (iwillrektyou):

The answer is either b or c.

OpenStudy (anonymous):

I got c.

OpenStudy (iwillrektyou):

Ok great job!

OpenStudy (anonymous):

Thank you